### v1

@@ 4.2 Heterogeneous distributed-lag regressions
4.2.1 Setup
We now study how we can identify and estimate expectations of $B$ under Assumption 7. First, we show that when combined with the standard parallel trends assumption (Assumption 4), Assumption 7 leads to a particular case of the random coefficients model considered by Chamberlain (1992). To see this, first note that Assumption 4 is equivalent to
@@ 4.2.1 Setup
where $\Delta$ is the first-difference operator. Next, let $\Delta\bm{Y}=(\Delta Y_{L+2},\dots,\Delta Y_{T})$, $\Gamma=(\gamma_{L+2},\dots,\gamma_{T})^{\prime}$, $\bm{\varepsilon}=(\varepsilon_{L+1},\dots,\varepsilon_{T})^{\prime}$ and
$\bm{M}=\begin{pmatrix}\Delta D_{L+2}&\dots&\Delta D_{2}\\ \Delta D_{L+3}&\dots&\Delta D_{3}\\ \vdots&\vdots&\vdots\\ \Delta D_{T}&\dots&\Delta D_{T-L}\end{pmatrix}$.
@@
We now introduce an identifying assumption on the design, which we impose to identify expectations of $B$. For any matrix $A$, let $A^{+}$ denote its Moore-Penrose inverse, $I$ denote the identity matrix (without specifying its size in the absence of ambiguity) and let $\Pi(A):=I-AA^{+}$ be the orthogonal projector on the orthocomplement of the image of $A$.
@@
(Data restriction) $T\geq 2(L+1)$ and $E[\Pi(\bm{M})]$ is invertible.
This condition only involves the treatment variable and is thus similar to Assumption 3, used in Section 3 above. The first condition $T\geq 2(L+1)$ imposes that $\bm{M}$ has at least as many rows as columns. Intuitively, this restriction is needed because even absent any time trends and noise in (13), if $T<2(L+1)$, one would not be able to recover $B$ from the equation $\Delta\bm{Y}=\bm{M}\times B$.
The second condition deserves more discussion. It plays the same role as the condition $F-1+\ell\leq\overline{T}$ that we imposed in the previous section to identify $\text{AVSQ}_{\ell}$, but it is weaker in general. To see this, assume for instance that $L=1$ and $T=5$. Then, one can show that $E[\Pi(\bm{M})]$ is invertible if for instance $\mathcal{D}$, the support of $\bm{D}$, satisfies $\mathcal{D}=\{(0,1,0,1,0),(0,0,1,1,0)\}$. Then, averages of $\beta_{0}$ and $\beta_{1}$, the effects of the treatment and its first lag, are identified. On the other hand, with such a design, $\text{AVSQ}_{2}$ is not defined because $P(D_{1}=D_{2}=D_{3})=0$. More generally, when $T>2(L+1)$, we can expect $E[\Pi(\bm{M})]$ to be invertible without requiring $P(D_{1}=...=D_{L+2})>0$, which is necessary, though not sufficient, to define $\text{AVSQ}_{L+1}$.
@@
When $T=2(L+1)$, $E[\Pi(\bm{M})]$ may not be invertible without stayers, namely groups that keep the same treatment over pairs of consecutive periods. If for instance the distribution of $\bm{D}$ is continuous with respect to the Lebesgue measure on $\mathbb{R}^{T}$, $P(\bm{M}\text{ is invertible})=1$ and thus $\Pi(\bm{M})=0$ a.s. Even with a binary treatment, if $L=1$ and $T=4$, one can show that a necessary condition for invertibility is
@@
so we need to have stayers over each triple of consecutive periods. That said, Assumption 8 can still hold if the treatment is continuously distributed across groups at each period, even if $T=2(L+1)$, because we may still have stayers. For instance, at any point in time a tax rate may be continuously distributed across regions, but that same tax rate may remain constant over time for several periods in many regions (de Chaisemartin et al., 2022).
4.2.2 Identification
Because $\Pi(\bm{M})\bm{M}=0$, it follows from (13) that under Assumption 8, $\Gamma$ is identified by
$\Gamma=E[\Pi(\bm{M})]^{-1}E[\Pi(\bm{M})\Delta\bm{Y}].$ (15)
Then, informally, we can learn about $\overline{B}:=E\left[B\big{|}\,\lambda_{\min}(\bm{M}^{\prime}\bm{M})>0\right]$ (with $\lambda_{\min}(A)$ the smallest eigenvalue of $A$) using
$\overline{B}=E\left[(\bm{M}^{\prime}\bm{M})^{-1}\bm{M}^{\prime}(\Delta\bm{Y}-\Gamma)\big{|}\lambda_{\min}(\bm{M}^{\prime}\bm{M})>0\right].$ (16)
The condition $\lambda_{\min}(\bm{M}^{\prime}\bm{M})>0$ states that there is enough independent variation in the treatment and its lags over time. In particular, $\lambda_{\min}(\bm{M}^{\prime}\bm{M})>0$ fails if $\Delta D_{t}=...=\Delta D_{t+k}=0$ for some $t\geq 2$ and $k\in\{T-L-2,\dots,T-t\}$. In the previous example with $L=1$, $T=4$ and $\mathcal{D}=\{(0,0,0,1),(1,0,1,0),(1,0,1,1)\}$, $\lambda_{\min}(\bm{M}^{\prime}\bm{M})>0$ fails when $\bm{D}=(0,0,0,1)$. It also fails if $\bm{D}=(1,0,1,0)$, because then $\Delta D_{t}=-\Delta D_{t-1}$ for both $t=3$ and $t=4$. On the other hand, the condition holds when $\bm{D}=(1,0,1,1)$.
Now, the identification argument in (16) is informal because as discussed in Graham and Powell (2012) and de Chaisemartin et al. (2022) in related setups, the expectation on the right-hand side of (16) may fail to exist. Namely, we may have
$E\left[\|(\bm{M}^{\prime}\bm{M})^{-1}\bm{M}^{\prime}(\Delta\bm{Y}-\Gamma)\|\,\big{|}\,\lambda_{\min}(\bm{M}^{\prime}\bm{M})>0\right]=\infty,$ (17)
due to “quasi-stayers”, that is to say, groups for which eigenvalues of $\bm{M}^{\prime}\bm{M}$ are arbitrary close to 0. In Appendix A.4, we show that indeed, (17) holds if $T=2$, $L=0$, $\Delta D_{2}$ has a positive density around 0 and a weak regularity condition holds on $\varepsilon_{2}$. Nevertheless, the following theorem shows that (i) we can still identify $\overline{B}$ by a suitable truncation and a continuity argument; (ii) this concern does not apply if $D_{t}$ is finitely supported for all $t$.
Theorem 4
Suppose that Assumption 1-4 and 7-8 hold and $E[\Delta D_{t}^{2}+\Delta Y_{t}(\bm{0}_{t})^{2}]<\infty$ for all $t\geq 2$. Then $\Gamma$ and $\overline{B}$ are identified respectively by (15) and
$\overline{B}=\lim_{\lambda\downarrow 0}E\left[(\bm{M}^{\prime}\bm{M})^{-1}\bm{M}^{\prime}(\Delta\bm{Y}-\Gamma)\big{|}\,\lambda_{\min}(\bm{M}^{\prime}\bm{M})>\lambda\right].$ (18)
Moreover, if $D_{t}$ is finitely supported for all t=1,,Tt=1,...,T, the right-hand side of (16) is well-defined and $\overline{B}$ is identified by (16).
Noteworthy, (18) does not rely on any regularity requirements such as continuity of some conditional expectations. This is because contrary to Graham and Powell (2012) and Sasaki and Ura (2021), we do not seek to extrapolate to groups for which $\lambda_{\min}(\bm{M}^{\prime}\bm{M})=0$.
The same reasoning as above shows that we can also identify
$\overline{B}(\mathcal{E}):=E\left[B\big{|}\,\bm{D}\in\mathcal{E},\,\lambda_{\min}(\bm{M}^{\prime}\bm{M})>0\right],$
for any set $\mathcal{E}\subseteq\mathcal{D}$. In particular, estimating $\overline{B}(\mathcal{E})$ may be helpful in an ex-ante perspective, to identify which groups should be treated in the future.
@@ 4.2.3 Estimation
Theorem 4 result suggests the following simple, plug-in estimators of $\Gamma$:
@@
This estimator is root-$G$ consistent and asymptotically normal, even when $T=2(L+1)$, as soon as Assumption 8 holds. This contrasts with Proposition 1.1 in Graham and Powell (2012), who consider a closely related random coefficients model: their result indicates that the common parameters ($\Gamma$ in our context, $\bm{\delta}_{0}$ in theirs) cannot be estimated at the standard, root-$G$ rate when the number of time periods is equal to the number of regressors ($T=2(L+1)$ in our context). An important difference with their setup is that we restrict the random coefficients to be time-invariant. This allows us to identify $\Gamma$ using observations for which $\bm{M}$ is singular.
Following Theorem 4, we discuss separately the estimation of $\overline{B}$ depending on whether $D_{t}$ is finitely supported or not. Note that estimation of $\overline{B}(\mathcal{E})$ follows along similar lines.
@@
5 Application to Gentzkow, Shapiro, and Sinkinson (2011)
@@
In this application, we revisit the effect of newspapers on voters’ participation in elections, studied by Gentzkow, Shapiro, and Sinkinson (2011). They use a US panel data set at the county $\times$ presidential-election level, with 1,195 counties and from the 1868 to the 1928 election. They seek to test a conjecture in De Tocqueville (1850), that newspapers encourage citizens to participate more in democratic institutions. For that purpose, they let $Y_{g,t}$ denote the turnout rate in county $g$ and the presidential election that took place in year $t$, they let $D_{g,t}$ denote the number of newspapers circulating in county $g$ and year $t$, and they run a static first-difference regression of $\Delta Y_{g,t}$ on $\Delta D_{g,t}$, and state-year fixed effects. Hereafter, we consider possibly dynamic effects. We first estimate the AVSQ event-study effects, under parallel trends only. Then, we estimate a standard distributed-lag TWFE regressions, before estimating the random coefficients distributed-lag linear model.
@@
First, note that the design of this application is truly complex: the number of newspapers is a non-binary treatment, it can increase or decrease over time, counties can experience several changes in their number of newspapers, at different points in time. The right panel of Figure 1 displays the distribution of the number of switches in $D_{1}$. It shows that over the sixteen years of presidential elections, counties experiencing eight or more changes are not uncommon. Actually, only 34 counties (2.9%) never experience any change in their number of newspaper. For 90.9% of the 1,161 remaining counties, $S_{g}=1$, namely they experience an increase in the number of newspapers at first switch. However, 78.7% of counties experience at least one decrease in their number of counties. Still, the no-crossing condition holds for 93.6% of the counties. Finally, the left panel of Figure 1 displays the distribution of $D_{1}$, which also shows significant heterogeneity. Assumption 3 is easily met, with $\mathcal{D}^{\text{r}}_{1}=\{0,1,...,7\}$.
Figure 1: Distribution of $D_{1}$ and the number of switches in $D_{1}$

### v3

@@ 4.2 Heterogeneous distributed-lag regressions
4.2.1 Motivation and setup
Under Assumption 7, the slope $\text{SL}^{\ell}_{k}$ defined in (7) reduces to $\beta_{k}$. Then, Theorem 2 shows that $\text{AVSQ}^{n}_{\ell}$ identifies a weighted average of expectations of the $\beta_{k}$s, for $k$ ranging from 0 to $\ell-1$. Actually, as discussed in a working paper version of de Chaisemartin and D’Haultfœuille (2025), under Assumption 7 one can leverage DID estimators to separately estimate the expectation of each $\beta_{k}$. First, one estimates group-specific regressions of
$(Y_{g,F_{g}-1+\ell}-Y_{g,F_{g}-1}-\widehat{\Delta}_{g,\ell})_{\ell\in\{1,...,\overline{T}-(F_{g}-1)\}},$
the building-block DID estimators in the previous section, on $(D_{g,F_{g}-1+\ell}-D_{g,1})_{\ell\in\{1,...,\overline{T}-(F_{g}-1)\}}$ and its lags. Then, one averages those group-specific regression coefficients across switchers (see Section 4 of de Chaisemartin and D’Haultfœuille, 2021). In this paper, we will not use DIDs to estimate the expectation of each $\beta_{k}$. Instead, we will propose estimators that may be used in more general designs, namely in designs that may not have stayers.
First, we show that when combined with the standard parallel trends assumption (Assumption 4), Assumption 7 leads to a particular case of the random coefficients model considered by Chamberlain (1992) and Arellano and Bonhomme (2012). To see this, first note that Assumption 4 is equivalent to
@@ 4.2.1 Motivation and setup
where $\Delta$ is the first-difference operator. Next, let $\Delta\bm{Y}=(\Delta Y_{K+2},\dots,\Delta Y_{T})$, $\Gamma=(\gamma_{K+2},\dots,\gamma_{T})^{\prime}$, $\bm{\varepsilon}=(\varepsilon_{K+1},\dots,\varepsilon_{T})^{\prime}$ and
$\bm{M}=\begin{pmatrix}\Delta D_{K+2}&\dots&\Delta D_{2}\\ \Delta D_{K+3}&\dots&\Delta D_{3}\\ \vdots&\vdots&\vdots\\ \Delta D_{T}&\dots&\Delta D_{T-K}\end{pmatrix}$.
@@
We now introduce an identifying assumption on the design, discussed in details below. For any matrix $A$, let $A^{+}$ denote its Moore-Penrose inverse, $I$ denote the identity matrix (without specifying its size in the absence of ambiguity) and let $\Pi(A):=I-AA^{+}$ be the orthogonal projector on the orthocomplement of the image of $A$.
@@
(Condition on the design) $E[\Pi(\bm{M})]$ is invertible.
4.2.2 Identification
We consider identification of average effects of the kind $E[\beta_{k}|A]$ for some subpopulation $A$ specified below. First, because $\Pi(\bm{M})\bm{M}=0$, it follows from (14) that under Assumption 8, $\Gamma$ is identified by
$\Gamma=E[\Pi(\bm{M})]^{-1}E[\Pi(\bm{M})\Delta\bm{Y}].$ (15)
Then, let $e_{k}$ denote the $k+1$-th canonical vector in $\mathbb{R}^{K+1}$ (e.g., $e_{0}=(1,0,...,0)^{\prime}$). Informally, we can identify $\overline{\beta}_{k}:=E[\beta_{k}|e_{k}\in\text{Im}(\bm{M}^{\prime})]$, provided that $P(e_{k}\in\text{Im}(\bm{M}^{\prime}))>0$, by
$\overline{\beta}_{k}=E\left[e^{\prime}_{k}\bm{M}^{+}(\Delta Y-\Gamma)|e_{k}\in\text{Im}(\bm{M}^{\prime})\right].$ (16)
Heuristically, (16) follows from the fact that when $e_{k}\in\text{Im}(\bm{M}^{\prime})$, $e^{\prime}_{k}\bm{M}^{+}\bm{M}=e^{\prime}_{k}$ and thus $e^{\prime}_{k}\bm{M}^{+}(\Delta Y-\Gamma)=\beta_{k}+e_{k}^{\prime}\bm{M}^{+}\bm{\varepsilon}$. This identification argument is informal, though, because as discussed in Graham and Powell (2012) and de Chaisemartin et al. (2022) in related setups, the expectation on the right-hand side of (16) may fail to exist. Namely, we may have
$E\left[|e^{\prime}_{k}\bm{M}^{+}(\Delta Y-\Gamma)|\,\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime})\right]=\infty,$ (17)
due to groups for which $\|\bm{M}^{\prime}{}^{+}e_{k}\|$ is arbitrary large. In Appendix A.4, we show that indeed, (17) holds if $T=2$, $K=0$, $\Delta D_{2}$ has a positive density around 0 and a weak regularity condition holds on $\varepsilon_{2}$. Nevertheless, the following theorem shows that (i) we can still identify $\overline{\beta}_{k}$ by a suitable truncation and a continuity argument; (ii) the concern above does not apply if $D_{t}$ is finitely supported for all $t$.
Theorem 4
Suppose that Assumptions 1, 2, 4, 7 and 8 hold and $E[\Delta D_{t}^{2}+\Delta Y_{t}(\bm{0}_{t})^{2}]<\infty$ for all $t\geq 2$. Then $\Gamma$ is identified by (15). Moreover, if $P(e_{k}\in\text{Im}(\bm{M}^{\prime})>0$ for $k\in\{0,...,K\}$, $\overline{\beta}_{k}$ is identified by
$\overline{\beta}_{k}=\lim_{C\uparrow\infty}E\left[e^{\prime}_{k}\bm{M}^{+}(\Delta Y-\Gamma)|e_{k}\in\text{Im}(\bm{M}^{\prime}),\|\bm{M}^{\prime}{}^{+}e_{k}\|\leq C\right]$ (18)
Finally, if $D_{t}$ is finitely supported for all t=1,,Tt=1,...,T, the right-hand side of (16) is well-defined and $\overline{\beta}_{k}$ is identified by (16).
This theorem is similar to Proposition 1 in Arellano and Bonhomme (2012), with a few differences. First, Model (14) is a particular case of their random coefficient model. Second, we consider the identification of averages of each component of $B$, rather than the identification of an average of the full vector $B$. This may be important in practice, because in the former case, we just need to restrict the estimation to groups for which $e_{k}\in\text{Im}(\bm{M}^{\prime})$, whereas in the latter case we need to restrict the estimation to groups for which $\bm{M}^{\prime}\bm{M}$ is invertible, and such groups may fail to exist (this is the case for instance if $T<2(K+1)$). The third difference with Proposition 1 in Arellano and Bonhomme (2012) is that we highlight the fact that identification can be achieved without trimming when $D_{t}$ is finitely supported, but may require trimming otherwise.
With respect to the identification results in Section 3, Theorem 4 applies to a broader class of designs. In particular, expectations of $(\beta_{k})_{k\in\{0,...,\ell-1\}}$ may be identified in designs that do not have stayers for long enough for $\text{AVSQ}_{\ell}$ to be identified, especially when $T>2(K+1)$, namely when the $(T-(K+1))\times(K+1)$ matrix $\bm{M}$ has more rows than columns. [^7] Intuitively, this helps for Assumption 8 because then, the image of $\bm{M}$ has dimension at most $K+1$, and thus the matrix $\Pi(\bm{M})$ has rank at least T2(K+1)>0T-2(K+1)>0. For instance, assume that $K=1$, $T=5$ and $\mathcal{D}$, the support of $\bm{D}$, satisfies $\mathcal{D}=\{(0,1,0,1,0),(0,0,1,1,0)\}$. Then, one can show that Assumption 8 holds, and $\overline{\beta}_{0}$ and $\overline{\beta}_{1}$ are identified. On the other hand, with such a design, $\text{AVSQ}_{2}$ is not identified because $P(D_{1}=D_{2}=D_{3})=0$. More generally, when $T>2(K+1)$ we can expect $E[\Pi(\bm{M})]$ to be invertible without requiring $P(D_{1}=...=D_{K+2})>0$, which is necessary to identify $\text{AVSQ}_{K+1}$. In fact, simulations suggest that Assumption 8 holds if $\bm{D}$ is continuous with respect to the Lebesgue measure on $\mathbb{R}^{T}$, a case where there is no stayer as $D_{2}\neq D_{1}$ and $F=2$ a.s, thus implying that Assumption 3 fails.
@@ 4.2.2 Identification
When $T\leq 2(K+1)$, $E[\Pi(\bm{M})]$ may not be invertible without groups that keep the same treatment value over consecutive periods, so averages of the random coefficients $(\beta_{k})_{k\in\{0,...,\ell-1\}}$ are identified under conditions similar to those under which $\text{AVSQ}_{\ell}$ is identified. For instance, Assumption 8 fails if $\bm{D}$ is continuous with respect to the Lebesgue measure on $\mathbb{R}^{T}$. Then, $P(\bm{M}\text{ is invertible})=1$ and $\Pi(\bm{M})=0$ a.s. Even with a binary treatment, if $K=1$ and $T=4$, one can show that a necessary condition for invertibility is
@@
a condition under which $\text{AVSQ}_{2}$ is typically identified.
@@ 4.2.3 Estimation
Theorem 4 suggests the following simple, plug-in estimator of $\Gamma$:
@@
This estimator is root-$G$ consistent and asymptotically normal, even when $T=2(K+1)$, as soon as Assumption 8 holds. This contrasts with Proposition 1.1 in Graham and Powell (2012), who consider a closely related random coefficients model: their result indicates that the common parameters ($\Gamma$ in our context, $\bm{\delta}_{0}$ in theirs) cannot be estimated at the standard, root-$G$ rate when the number of time periods is equal to the number of regressors ($T=2(K+1)$ in our context). An important difference with their setup is that we restrict the random coefficients to be time-invariant. This allows us to identify $\Gamma$ using observations for which $\bm{M}$ is singular.
Following Theorem 4, we discuss separately the estimation of $\overline{\beta}_{k}$ depending on whether $D_{t}$ is finitely supported or not.
@@
4.3 Further results
4.3.1 Allowing for dynamic effects up to any treatment lag
In this section, we replace Assumption 7 by the following condition:
Assumption 9
(Random coefficients distributed-lag linear model, with effects up to any lag) There exists a random vector $B:=(\beta_{0},...,\beta_{T-1})^{\prime}$ such that $E[\|B\|^{2}]<\infty$ and
$Y_{t}\left(d_{1},...,d_{t}\right)=Y_{t}(\bm{0}_{t})+\sum_{k=0}^{t-1}\beta_{k}d_{t-k}.$
Then, we have
$\Delta Y=\Gamma+\bm{M}B+\bm{\varepsilon},$
where now $\Delta Y=(\Delta Y_{2},...,\Delta Y_{T})^{\prime}$, $\bm{\varepsilon}=(\varepsilon_{2},...,\varepsilon_{T})^{\prime}$ $\Gamma=(\gamma_{2},...,\gamma_{T})^{\prime}$ and
$\bm{M}=\begin{pmatrix}\Delta D_{2}&0&&\dots&0\\ \Delta D_{3}&\Delta D_{2}&0&\dots&0\\ \vdots&&&\ddots&\vdots\\ \Delta D_{T}&&\dots&&\Delta D_{2}\end{pmatrix}$.
Then, identification and estimation under Assumption 9 proceeds as in Sections 4.2.2 and 4.2.3. Remark that $\bm{M}$ is now a $(T-1)\times(T-1)$ matrix. Then, one can show that Assumption 8 holds if and only if there are some “never switchers”, namely $P(D_{1}=...=D_{T})>0$.
Assumption 9 allows all lagged treatments up to period one to affect the outcome. Thus, on that dimension it is less restrictive than Assumption 7, which only allows the first $K$ lags to have an effect. At the same time, note that under Assumption 9,
$Y_{t}=Y_{t}(\bm{0}_{t})+\sum_{k=0}^{t-1}\beta_{k}D_{t-k}:$ (21)
the number of lags allowed to affect the outcome depends on $t$, the period when the outcome is measured. This reflects an implicit assumption embedded in our potential outcome notation, namely that treatments prior to period one have no effect on groups’ outcomes from period one to $T$. If all lags can affect the outcome, namely if instead of (21) we have that
$Y_{t}=Y_{t}(\bm{0}_{t})+\sum_{k=0}^{+\infty}\beta_{k}D_{t-k},$
then the estimators introduced in this section can still be used if groups’ treatments do not change before period 1:
$\forall t\leq 0,\penalty 10000\ D_{t}=D_{1}\text{ almost surely.}$ (22)
(22) is a strong condition. Still, it holds by construction when the treatment does not exist before period one. When the treatment is absorbing, it also holds by construction in the subsample of untreated groups at period one. If treatments prior to period one have an effect on groups’ outcomes and groups’ treatments changed before period one, then the estimators introduced in this section may be misleading. Instead, the estimators proposed in Section 4.2.3 can be used in such cases, at the expense of ruling out effects of lagged treatments beyond the $K$th lag, and dropping the outcomes $(Y_{t})_{t=1,...,K}$ from the estimation.
4.3.2 Allowing for interaction effects
In this section, we replace Assumption 7 by the following condition, which allows for interaction effects between treatment lags:
Assumption 10
(Random coefficients distributed-lag linear model with interactions) There exist random variables $(\beta_{k,k^{\prime}})_{0\leq k\leq k^{\prime}\leq K}$ such that
$Y_{t}(d_{1},...,d_{t})=Y_{t}(\bm{0}_{t})+\sum_{k=0}^{K}\left[\beta_{k,k}+\sum_{k^{\prime}=k+1}^{K}\beta_{k,k^{\prime}}d_{t-k^{\prime}}\right]d_{t-k},$
where K<T1K<T-1 is known. Moreover, $E[\beta_{k,k^{\prime}}^{2}]<\infty$ for all $0\leq k\leq k^{\prime}\leq K$.
Then, (14) still holds, except that now the matrix $\bm{M}$ is of dimension $(T-K+1)\times\frac{(K+1)(K+2)}{2}$, and includes columns of the kind $\Delta(D_{t-k}D_{t-k^{\prime}})$ for some fixed $(k,k^{\prime})$ and where $t$ varies over the column. As a result, identification may fail even in some canonical designs. For instance, with a binary and absorbing treatment, one cannot tease out the effects of $D_{t-1}D_{t}$ and $D_{t-1}$: as no treated group goes back to being untreated, $D_{t-1}D_{t}=D_{t-1}$. [^8] Formally, the identifying condition corresponding to $P(e_{k}\in\text{Im}(\bm{M}^{\prime}))>0$ in Theorem 4 does not hold. On the other hand, identification may still be achieved with non-absorbing treatments. For instance, if $K=1$, $T=5$ and $\mathcal{D}=\{(0,0,0,0,0),(0,1,1,0,0)\}$, we can identify averages of $\beta_{0,0}$, $\beta_{0,1}$, and $\beta_{1,1}$ on groups for which $\bm{D}=(0,1,1,0,0)$.
5 Application to Gentzkow et al. (2011)
@@ 4.3.2 Allowing for interaction effects
In this application, we revisit the effect of newspapers on voters’ participation in elections, studied by Gentzkow et al. (2011). They use a US panel data set at the county $\times$ presidential-election level, with 1,195 counties and from the 1868 to the 1928 election. They seek to test a conjecture in De Tocqueville (1850), that newspapers encourage citizens to participate more in democratic institutions. For that purpose, they let $Y_{g,t}$ denote the turnout rate in county $g$ and the presidential election that took place in year $t$, they let $D_{g,t}$ denote the number of newspapers circulating in county $g$ and year $t$, and they run a static first-difference regression of $\Delta Y_{g,t}$ on $\Delta D_{g,t}$, and state-year fixed effects. Hereafter, we consider possibly dynamic effects. We first estimate the AVSQ event-study effects, under parallel trends only. Then, we estimate a standard distributed-lag TWFE regressions, before estimating the random coefficients distributed-lag linear model.
@@
First, note that the design of this application is truly complex: the number of newspapers is a non-binary treatment, it can increase or decrease over time, counties can experience several changes in their number of newspapers, at different points in time. The right panel of Figure 1 displays the distribution of the number of switches in $D_{t}$. It shows that over the sixteen years of presidential elections, counties experiencing eight or more changes are not uncommon. Actually, only 34 counties (2.9%) never experience any change in their number of newspapers. For 90.9% of the 1,161 remaining counties, $S_{g}=1$, namely they experience an increase in the number of newspapers at first switch. However, 78.7% of counties experience at least one decrease in their number of newspapers. Still, the no-crossing condition holds for 93.6% of the counties. Finally, the left panel of Figure 1 displays the distribution of $D_{1}$, which also shows significant heterogeneity. Assumption 3 is easily met, with $\mathcal{D}^{\text{r}}_{1}=\{0,1,...,7\}$.
Figure 1: Distribution of $D_{1}$ and the number of switches in $D_{t}$
Notes: the last bins (5 and 10) actually correspond to 5 or more and 10 or more.